OpenStudy (anonymous):

Given that (-3,-5) and (-1,-4) are points on the same line, find a point-slope form of the equation for this line??

pooja195 (pooja195):

Hi welcome to openstudy :) Start by finding the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

\[\LARGE \frac{-4--5}{-1--3}=slope\] \[\LARGE \frac{-4+5}{-1+3}=slope\]